The length of a rectangle is twice the width. The perimeter is 48 in. Find the dimensions of the rectangle?
Answer by elizabeth2005(20)


what you do is write out your problem:
the length of the rectangle is twice the width:
so..2(w)
then you have to add the width to that problem
2(w)+w=48
you now multiply 2 by everything in the par. wich leaves you with 2w+w=48
now you have to combine like terms...2w+w=3w...
3w=48 divide both sides by 3
three divided by 48 leaves you with 16
next you have to multiply 2 by 16 and it leaves you with 32
your length should be 32
and the width should be 16
if you want you can add 32 to 16 and you will get 48